OpenStudy (anonymous):

log28 + log22 expressed as a single logarithm is: log416 log410 log216 log210

OpenStudy (raden):

log a + log b = log (a * b)

OpenStudy (anonymous):

so does that include the numbers under the log

OpenStudy (raden):

2 as the base, right ? log2 8 + log2 2 = log2 (8 * 2) = log2 16

OpenStudy (anonymous):

oh okay thanks

OpenStudy (raden):

welcome